The chapter demonstrated that a firm borrowing in a foreign currency could potentially end up paying a very different effective rate of interest than what it expected. Using the same baseline values of a debt principal of SF1.5 million, a​ one-year period, an initial spot rate of SF1.4600​/$, a 5.302​% cost of​ debt, and a 38​% tax​ rate, what is the effective​ after-tax cost of debt for one year for a U.S.​ dollar-based company if the exchange rate at the end of the period​ was:

a. SF1.4600​/$

b. SF1.4100​/$

c. SF1.3670​/$

d. SF1.5840​/$

Round to four decimal places
Answer #1

Debt principal(Swiss francs)=1,500,000

Initial spot rate(s1)=1.4600

Cost of debt=5.302%

Tax rate=38.00%

a. If the exchange remains the same= 1.4600

We first calculate the percentage change in the exchange rate, s = ( S1 - S2 ) / (S2)x100 =1.46-1.46 /1.46 x100=0.000%

We then calculate the effective cost of debt after exchange rate changes=5.302%

b. If the exchange ends the period at SF1.4100/$:

We first calculate the percentage change in the exchange rate, s = ( S1 - S2 ) / (S2)x100 =1.46-1.41 /1.41 x100=3.5461%

We then calculate the effective cost of debt after exchange rate changes

kd= [ ( 1 + kd in SF) x ( 1 + s ) ]- 1 =[(1+.05302)x(1+0.03546)]-1=9.0360%

c. If the exchange ends the period at SF1.3670/$:

We first calculate the percentage change in the exchange rate, s = ( S1 - S2 ) / (S2)x100=6.8032%

We then calculate the effective cost of debt after exchange rate changes

kd= [ ( 1 + kd in SF) x ( 1 + s ) ]- 1 =[(1+.05392)x(1+0.068032)]-1=12.5620%.

d.

If the exchange ends the period at SF1.5840/$:

We first calculate the percentage change in the exchange rate, s = ( S1 - S2 ) / (S2)x100=-7.8283%

We then calculate the effective cost of debt after exchange rate changes

kd= [ ( 1 + kd in SF) x ( 1 + s ) ]- 1

=[(1+.05302)x(1-0.078283)]-1=-2.9413%